A pure titanium cube has an edge length of 2.84in

How many titanium atoms does it contain? Titanium has a density of 4.50g/cm3.
Answer #1

x = multiply
/ = divide (fraction)
^ = power, exponent


Volume of titanium cube:
2.84 x 2.84 x 2.84 (or 2.84^3) = 22.906 cubic ...cm? you didn't specify units. if it's not CM you'll have to do some conversions.

Density of Ti: 4.506 g/cm3

4.506g/cm3 x 22.906 cm3 = 103.214 grams

103.214 grams / 47.867 grams/mol = 2.156 mols

2.156 mols x 6.023x10^23 molecules/mol = 1.2986x10^24 molecules

should also mean 1.2986x10^24 atoms?
Answer #2

Correction:

Mass of the cube:
M = (2.84 * 25.4 * 10?3 m)